# Evaluating Task-based Effectiveness of MLLMs on Charts

### v1

@@ 4.1.1 Overall Evaluation on Various MLLMs
Reasoning
Anomaly
Distribution
Correlation
Range
Order
Filter
Retrieval
Extremum
Cluster
Open Source MLLMs
VisCPM-Chat-v1.1  [35] 28.4 46.1 33.3 51.9 23.0 6.4 25.1 15.8 32.0 29.6 26.2
BLIP2  [36] 24.8 23.4 25.0 15.1 25.3 20.2 39.8 27.8 30.3 30.1 28.3
CogVLM-17B  [37] 20.3 23.1 43.6 29.6 37.7 10.8 9.1 37.9 56.6 26.7 29.4
OmniLMM-12B  [38] 24.7 19.9 27.0 34.9 35.7 28.3 30.0 33.0 39.9 33.1 31.1
LLaVA1.5  [39] 32.4 6.3 30.9 23.1 21.7 32.7 35.6 32.6 35.8 43.5 32.2
ChartAssistant  [40] 24.6 27.7 35.8 28.1 30.5 22.5 14.7 39.4 63.0 26.4 32.4
MiniCPM-v2  [41] 19.5 55.1 33.3 56.5 24.9 16.7 36.3 37.9 52.4 32.0 33.0
mPLUG-Owl2  [42] 31.0 27.0 29.4 35.3 28.4 22.5 40.3 30.9 41.1 27.3 33.3
Qwen-VL-Chat  [43] 27.8 36.3 45.1 55.8 33.8 20.0 28.7 31.3 50.2 27.1 33.4
ViP-LLaVA  [44] 28.8 6.6 34.8 30.3 21.9 35.8 40.4 42.2 38.3 33.8 33.8
LLaVA-NEXT  [45] 30.6 7.4 26.5 38.0 29.5 33.3 23.4 53.5 59.8 52.3 38.5
Sphinx-v2  [46] 30.0 28.9 37.8 36.1 25.8 23.5 36.7 49.7 66.3 45.3 40.2
Closed Source MLLMs
Qwen-VL-Plus [47] 30.8 27.3 47.1 47.1 43.0 34.6 20.7 58.7 65.5 62.5 42.6
Gemini-Pro-Vision  [28] 25.6 30.1 45.6 58.7 75.3 32.9 30.1 60.4 80.9 55.3 48.4
ChatGLM-4V  [48] 34.1 28.9 39.2 42.3 55.5 18.9 43.4 58.1 69.3 71.4 48.4
Claude3-Haiku  [49] 33.0 9.0 42.7 46.2 60.4 26.2 40.0 62.3 75.1 66.8 49.5
Qwen-VL-Max [47] 28.8 25.8 62.3 63.0 66.1 40.2 38.9 67.0 79.6 66.8 51.7
GPT-4V  [50] 35.2 19.5 53.4 59.6 70.0 41.9 44.3 67.6 88.7 72.9 56.1
Table 3: Accuracy scores of 18 MLLMs across 7 chart and 4 question types on ChartInsights. FB: Fill-in-the-Blank Prompt; MC: Multiple-Choice Prompt; YN: Yes-or-No Prompt; EC: Error Correction Prompt.
@@
VisCPM-Chat-v1.1  [35] 24.8 21.5 24.4 30.3 25.3 34.9 20.4 7.4 39.8 49.3 8.3
BLIP2  [36] 32.2 31.1 24.7 34.2 15.0 18.2 8.7 3.1 46.9 57.7 5.5
CogVLM-17B  [37] 26.2 24.2 32.1 40.1 35.3 30.2 19.6 27.6 49.7 31.3 9.2
OmniLMM-12B  [38] 28.8 25.9 26.5 37.8 43.4 27.5 34.4 10.1 46.3 57.0 10.8
LLaVA1.5  [39] 31.3 30.5 22.6 35.7 40.6 26.8 36.3 9.6 41.4 67.9 9.9
ChartAssistant  [40] 32.9 27.8 24.1 41.2 35.6 28.2 23.3 30.2 46.3 40.9 12.2
MiniCPM-v2  [41] 33.4 28.7 18.8 40.4 36.3 29.9 28.0 21.0 42.0 60.8 8.3
mPLUG-Owl2  [42] 32.4 31.8 29.2 37.0 38.4 29.2 34.3 13.2 49.9 54.6 15.5
Qwen-VL-Chat  [43] 31.2 26.3 25.9 39.3 42.8 38.5 34.7 22.5 52.3 48.6 10.4
ViP-LLaVA  [44] 32.2 31.2 16.1 39.8 34.4 28.1 39.1 7.9 40.1 73.4 13.7
LLaVA-NEXT  [45] 37.1 33.7 18.8 50.4 48.4 27.4 37.1 23.1 42.2 63.2 25.6
Sphinx-v2  [46] 39.4 35.7 26.5 51.2 42.5 31.5 36.7 28.2 56.1 60.0 16.7
@@
Qwen-VL-Plus [47] 36.5 35.8 26.2 57.0 32.8 44.2 42.5 32.9 59.2 54.0 24.2
Gemini-Pro-Vision  [28] 45.0 43.3 35.1 57.7 43.8 47.3 51.2 40.9 55.4 43.9 53.3
ChatGLM-4V  [48] 46.0 43.9 36.9 58.7 56.3 39.7 49.9 32.1 53.7 72.2 35.8
Claude3-Haiku  [49] 50.7 48.1 32.1 55.6 46.9 40.4 48.3 39.4 51.2 61.8 45.9
Qwen-VL-Max [47] 47.1 46.3 38.4 66.4 43.1 47.7 48.1 44.2 75.3 48.1 39.0
GPT-4V  [50] 52.0 48.2 47.3 67.2 49.4 62.7 53.6 44.1 64.4 66.4 49.7
4.1.2 In-depth Evaluation of Textual Prompt Variations on GPT-4V
In this set of experiments, we aim to achieve two main goals: First, to benchmark GPT-4V’s performance across 10 low-level ChartQA tasks. Second, to investigate the impact of 4 types of Textual Prompt strategies on GPT-4V.
Figure 6: The Overall Accuracy: Task Groups vs. Chart Types
Overall Results. Figure 6 displays GPT-4V’s performance across a range of chart types and task categories, highlighting a notably low overall accuracy within the Analysis task category. Specifically, GPT-4V shows its worst performance on stacked bar charts, with a mere average accuracy of 19.8%. Conversely, its strongest performance is observed in the Query task category, particularly with scatter plots, where it achieves an impressive accuracy of 89.8%.
The disparity in GPT-4V’s experimental results can be attributed primarily to the nature of the tasks within each category. The Analysis task category includes a range of data analysis tasks that require complex reasoning, calculations, determination of correlations, understanding of data distributions, and identification of anomalies. In contrast, the Query task category involves simpler tasks, such as acquiring specific data values, which are inherently less complex.
@@ 4.1.2 In-depth Evaluation of Textual Prompt Variations on GPT-4V
As depicted in the bar chart at the top of Figure 6, GPT-4V’s accuracy in low-level ChartQA tasks reaches just over 60% for basic bar and scatter plots, while it hovers around 50% for similar tasks involving other chart types. Despite the dataset containing a significant number of reasoning tasks, these tasks are generally straightforward for humans. This performance gap suggests that, in the context of these tests, GPT-4V’s current level of chart comprehension has not yet reached that of the average human.
Finding-2: The overall accuracy of GPT-4V gradually decreases as the difficulty of the task increases, which is similar to human performance on chart understanding tasks, but GPT-4V cannot reach a level of chart understanding and analysis similar to that of ordinary humans.
The Effectiveness of Textual Prompts. Table 4 presents the overall performance of GPT-4V across 10 low-level tasks with four textual prompts. Specifically, GPT-4V exhibits the highest overall accuracy with the Yes-or-No prompt, achieving 66.39%. In addition, it also performs well with the Multiple-Choice prompt, with 64.35% accuracy.
GPT-4V shows better performance on Multiple-Choice, Yes-or-No, and Error Correction prompts than the Fill-in-the-Blank prompt. The first three prompt types inherently offer candidate answers, allowing GPT-4V to select or judge, whereas Fill-in-the-Blank demands direct answer generation from GPT-4V.
@@
Finding-3: Structured textual prompts and candidate answers significantly enhance GPT-4V’s ability to reason out correct responses.
The Effectiveness of Chart Types. Table 5 shows the overall accuracy of GPT-4V for different chart types in 10 low-level tasks. Overall, GPT-4V has the highest performance on basic bar charts, reaching an average accuracy of 67.24%. The main reason is that the chart structure of the basic bar chart is relatively simple. Similarly, GPT-4V achieves better results on charts with simple structures such as scatter plots and pie charts. For charts with complex structures, such as stacked bar charts, grouped bar charts, and grouped line charts, the average accuracy of GPT-4V is lower than 50%. Specifically, GPT-4V achieved the worst performance in the three low-level tasks on the stacked bar chart, accounting for 50% (3/6).
Finding-4: The complexity of the chart structures and visual elements will significantly affect the performance of GPT-4V on low-level tasks.
Table 4: The Effectiveness of Textual Prompts vs. Ten Low-level Tasks
Textual Prompts Analysis Search Query Overall (%)
Reasoning
Anomaly
Distribution
Correlation
Range
Order
Filter
Retrieval
Extremum
Cluster
Fill-in-the-Blank 16.59 15.62 54.9 50 49.35 20 20.85 67.47 89.24 73.3 44.05
Multiple-Choice 35.62 17.19 78.43 64.42 99.22 43.08 48.14 81.89 97.85 82.39 64.35
Yes-or-No 66.46 31.25 21.57 72.12 54.57 78.46 75.42 55.45 79.8 67.61 66.39
Error Correction 21.98 14.06 58.82 51.92 77.02 26.15 32.88 65.54 87.75 68.18 49.73
Overall Accuracy (%) 35.17 19.53 53.43 59.62 70.04 41.92 44.32 67.59 88.66 72.87 56.13
Table 5: The Average Accuracy of Different Chart Types vs.Ten Low-level Tasks (“–” means “N/A”)
Chart Types Analysis Search Query Overall (%)
Reasoning
Anomaly
Distribution
Correlation
Range
Order
Filter
Retrieval
Extremum
Cluster
Grouped Bar 21.89 77.63 41.45 64.87 92.63 65 51.97
Stacked Bar 19.79 54.41 43.2 66.91 74.08 81.25 48.22
Grouped Line 22.02 71.43 73.81 47.32
Basic Bar 65.36 32.69 74.15 46.01 46.73 72.02 94.74 67.24
Basic Line 52.34 37.5 67.19 49.38
Scatter Plot 55.15 16.18 53.43 68.63 66.67 90.2 86.54 62.71
Pie 48.86 72.73 31.25 47.44 63.92 53.56
4.2 The Impact of Visual Variations and Visual Prompts
Most ChartQA evaluations [8, 9, 10, 11, 12] focus on the impact of Textual Prompt, failing to consider the chart’s quality and the Visual Prompts. Therefore, our research aims to explore how visual variations and visual prompts influence GPT-4V’s performance.
Figure 7: The Impact of Visual Variations. The heatmap shows how the performance of GPT-4V is affected by varying different chart elements.
Figure 8: The Impact of Image Quality. The heatmap shows how the performance of GPT-4V is affected by varying the image quality.
4.2.1 Varying Chart Elements
A visualization chart can exhibit visual differences by varying its elements, as shown in Figure 3(a). Intuitively, these visual differences are likely to influence GPT-4V’s performance on low-level ChartQA tasks. Thus, we explore how varying chart elements affect the performance of GPT-4V in this set of experiments.
Experimental Settings. The settings for varying chart elements is illustrated in Figure 3(a). As shown in Figure 2(a)-Step 4, we use the 356 visual variants for 35 charts as the testing samples. These 356 visual variants (charts) are associated with 17,972 textual prompts and cover 10 low-level tasks.
Overall Results. We calculate the overall accuracy of GPT-4V and compare it with the results in Tables 4 and 5 to record the change in performance. Figure 7 reports the experimental results.
Overall, the color of most areas of the heatmap in Figure 7(a) is light yellow, indicating that most chart variants have a slight negative impact on the performance of GPT-4V. In settings without data labels, GPT-4V’s performance declines significantly, particularly across seven types of charts (Figure 7(a)). This outcome is understandable, as data labels facilitate GPT-4V’s comprehension of the underlying insights conveyed by the charts. However, an interesting pattern emerges: as shown in Figure 7(b), GPT-4V shows a performance improvement of 17.5% in anomaly detection and 5.5% in filtering tasks when data labels are absent. This suggests that data labels might, in some instances, hinder GPT-4V’s ability to identify anomalies and filter values effectively.
As depicted in the heatmap of Figure 7(b), it is evident that several chart variants, such as larger x/y/data labels, positively affect GPT-4V’s performance in tasks like anomaly detection, filtering, ordering, and clustering. These tasks inherently involve comparisons between elements. We hypothesize that, in many cases, alterations in chart elements may shift GPT-4V’s focus towards visual comparisons rather than numerical ones, thereby enhancing its performance in these tasks.
The bar chart on the right in Figure 7 depicts the varied impacts of 15 chart variants on GPT-4V’s performance. We believe that data labels play a crucial role in GPT-4V’s low-level data analysis capabilities, as removing them or reducing their size tends to diminish its effectiveness. Moreover, adding marks to the legend or eliminating the legend’s color negatively affects GPT-4V by introducing visual clutter and removing essential visual cues, respectively.
Finding-5: While most chart variants slightly hinder GPT-4V’s performance, especially in the absence of data labels, certain modifications like larger labels and the removal of data labels can actually improve its performance in tasks like anomaly detection and filtering, by shifting its focus to visual comparisons.
4.2.2 Varying Image Quality
In addition to how visual modifications to chart elements alter the order and manner in which we interpret charts, the quality of chart images also plays a critical role in humans’ comprehension of these visual representations. Thus, it raises a compelling question: do these factors, known to have varying degrees of negative impact on human understanding, similarly impede GPT-4V’s ability to decipher charts?
Experimental Settings. The settings for varying chart elements is illustrated in Figure 3(b). We introduce six types of noise to evaluate the robustness and reliability of GPT-4V in low-level ChartQA tasks. As shown in Figure 2(a)-Step 4, we use the 245 visual variants for 35 charts as the testing samples. These 245 visual variants (charts) are associated with 8,456 textual prompts and cover 10 low-level tasks.
Overall Results. We calculate the overall accuracy of GPT-4V and compare it with the results in Tables 4 and 5 to record the change in performance. Figure 8 reports the experimental results. Generally, six methods of degrading image quality tend to negatively impact GPT-4V across a broad range of tasks and chart types. Among these, Median Blur stands out as the most detrimental, causing an average performance decline of 14.8%. We consider that median blurring makes numerical labels unreadable, resulting in a significant decrease in the performance of tasks directly related to numerical values.
Interestingly, adjustments to brightness—both increasing and decreasing—show a positive effect on the majority of tasks, with an average improvement of 0.6% and 1.4%, respectively.
As shown in Figure 8(b), the distribution task presents a unique case; it is adversely affected solely by Median Blur, whereas other forms of image quality manipulation tend to improve its performance.
Finding-6: The readability of the chart after median blurring is very poor and difficult to fully understand even for humans. Thus, it’s reasonable to assume that GPT-4V likewise faces difficulties in processing charts with compromised readability.
4.2.3 The Impact of Visual Prompts
In the field of computer vision, many studies aim to enhance models’ semantic extraction and entity recognition in images through the use of visual prompts [51]. Unlike general VQA (visual question answering) tasks, the ChartQA task, especially for our low-level data analysis tasks, demands greater sensitivity to detail and higher accuracy from the model. Therefore, we design different visual prompts for different low-level tasks to help GPT-4V better adapt to the requirements of different low-level tasks.
@@ 4.2.3 The Impact of Visual Prompts
Experimental Settings. We design three types of visual prompts, namely, hand writing, regular shape, and special design, as shown in Figure 4. As shown in Figure 2(a)-Step 5, we generate the 255 visual prompts for 35 charts as the testing samples. These 255 visual prompts are associated with 1,024 textual prompts and cover 10 low-level tasks.
@@
Overall Results. Figure 5(b) reports the experimental results. We can see that GPT-4V equipped with visual prompts shows strong performance across 10 distinct tasks and 4 textual prompts, demonstrating its effectiveness. As shown in Figure 9(a)-(c), we also show the performance of visual prompts under different textual prompts, low-level ChartQA tasks, and chart types. Generally, visual prompts have been shown to enhance GPT-4V’s performance. Notably, as depicted in Figure 9(b), GPT-4V equipped with visual prompts demonstrates great improvements in Reasoning and Anomaly Detection tasks. This indicates that visual prompts enable the model to accurately capture relevant data for analysis and reasoning.
Finding-7: Visual prompts significantly improve GPT-4V’s performance in various scenarios, such as textual prompts, low-level tasks, and chart types, highlighting the utility of visual information in aiding comprehension and reasoning.

### v3

@@ 3.1 Experimental Design
Table 2: Performance of 19 MLLMs across 10 Low-level ChartQA Tasks.
Table 3: Performance of 19 MLLMs across 7 Chart Types and 4 Question Types. (FB: Fill-in-the-Blank Question; MC: Multiple-Choice Question; YN: Yes-or-No Question; EC: Error Correction Question)
@@
VisCPM-Chat-v1.1  Hu et al. (2023a) 24.8 21.5 24.4 30.3 25.3 34.9 20.4 7.4 39.8 49.3 8.3
BLIP2  Li et al. (2023b) 32.2 31.1 24.7 34.2 15.0 18.2 8.7 3.1 46.9 57.7 5.5
CogVLM-17B  Wang et al. (2024) 26.2 24.2 32.1 40.1 35.3 30.2 19.6 27.6 49.7 31.3 9.2
OmniLMM-12B  Modelbest (2024) 28.8 25.9 26.5 37.8 43.4 27.5 34.4 10.1 46.3 57.0 10.8
LLaVA1.5  Liu et al. (2024b) 31.3 30.5 22.6 35.7 40.6 26.8 36.3 9.6 41.4 67.9 9.9
ChartAssistant  Meng et al. (2024) 32.9 27.8 24.1 41.2 35.6 28.2 23.3 30.2 46.3 40.9 12.2
MiniCPM-v2  Hu et al. (2024) 33.4 28.7 18.8 40.4 36.3 29.9 28.0 21.0 42.0 60.8 8.3
mPLUG-Owl2  Ye et al. (2023) 32.4 31.8 29.2 37.0 38.4 29.2 34.3 13.2 49.9 54.6 15.5
Qwen-VL-Chat  Bai et al. (2023) 31.2 26.3 25.9 39.3 42.8 38.5 34.7 22.5 52.3 48.6 10.4
ViP-LLaVA  Cai et al. (2023) 32.2 31.2 16.1 39.8 34.4 28.1 39.1 7.9 40.1 73.4 13.7
LLaVA-NEXT  Liu et al. (2024a) 37.1 33.7 18.8 50.4 48.4 27.4 37.1 23.1 42.2 63.2 25.6
Sphinx-v2  Lin et al. (2023) 39.4 35.7 26.5 51.2 42.5 31.5 36.7 28.2 56.1 60.0 16.7
@@
Qwen-VL-Plus Bai et al. (2023) 36.5 35.8 26.2 57.0 32.8 44.2 42.5 32.9 59.2 54.0 24.2
Gemini-Pro-Vision  Team et al. (2024) 45.0 43.3 35.1 57.7 43.8 47.3 51.2 40.9 55.4 43.9 53.3
ChatGLM-4V  Zeng et al. (2023) 46.0 43.9 36.9 58.7 56.3 39.7 49.9 32.1 53.7 72.2 35.8
Claude3-Haiku  Anthropic (2024) 50.7 48.1 32.1 55.6 46.9 40.4 48.3 39.4 51.2 61.8 45.9
Qwen-VL-Max Bai et al. (2023) 47.1 46.3 38.4 66.4 43.1 47.7 48.1 44.2 75.3 48.1 39.0
GPT-4V  OpenAI (2023) 52.0 48.2 47.3 67.2 49.4 62.7 53.6 44.1 64.4 66.4 49.7
GPT-4o  OpenAI et al. (2024) 62.7 59.8 53.9 84.1 56.9 74.6 70.9 61.0 77.8 74.8 63.1
3.2 Experimental Results and Findings
◆ Exp-1: Evaluation and Benchmarking
Experimental Settings.
We evaluate 19 widely used models (see Table 2) from both academia and industry, including 12 open-source and 7 closed-source MLLMs. Inspired by existing evaluation strategies Lu et al. (2024); Fu et al. (2024), we randomly selected 20% of the ChartInsights dataset as our test set to reduce testing costs while ensuring the reliability of experimental results.
The test set comprises 400 charts, spanning 10 low-level analysis tasks and encompassing 4,388 (chart, task, query, and answer) samples. This sizable validation dataset allows for a comprehensive evaluation of MLLMs’ capabilities in low-level ChartQA tasks. Then, we analyze the answers of MLLMs, compare them with Ground Truth, and calculate the accuracy. For each model, we evaluate through four question types in textual prompts, i.e., Fill-in-the-Blank, Multiple-Choice, Yes-or-No, and Error Correction questions.
Overall Results.
Table 2 presents the performance of the 19 models across 10 low-level tasks, while Table 3 showcases their performance across 7 chart types and 4 question types. The findings indicate that closed-source models outperform open-source models by a significant margin among the 19 models evaluated. Notably, GPT-4o demonstrates exceptional performance across all tasks.
Finding-1: Closed-Source models exhibit far superior generalization performance in low-level analysis tasks compared to open-source models.
In addition, our analysis of open-source models reveals that the ability to comprehend low-level ChartQA tasks is not directly proportional to the number of model parameters. For example, MiniCPM (with 2B parameters) outperforms larger models like CogVLM (with 17B parameters). This suggests that factors beyond scale significantly influence a model’s capacity.
Finding-2: The ability of open-source models to understand low-level charts is not directly proportional to their number of model parameters.
Moreover, Table 3 shows that although some open-source models perform poorly overall, they have high accuracy on Yes-or-No tasks, with ViP-LLaVA’s accuracy even being close to that of GPT-4o. We analyze the recall rates of ViP-LLaVA and GPT-4o on Yes-or-No tasks based on the confusion matrix and found that ViP-LLaVA’s recall rate for the “No” label is only 25%, which is quite different from the final accuracy of 73.4%; while GPT-4o’s recall rate for the “No” label is 72%, which is relatively close to the final accuracy of 74.8%. We believe that open-source models like ViP-LLaVA have a certain tendency towards the “No” label, and it is because there are more “No” labels in our data that these open-source models have a high accuracy on Yes-or-No questions.
Finding-3: We find open-source models like ViP-LLaVA show higher accuracy on Yes-or-No questions because of a possible bias towards “No” labels.
GPT-4o as a representative research object.
Since GPT-4o significantly outperforms other MLLMs, we utilize it as a representative MLLM to systematically evaluate the effectiveness of low-level ChartQA tasks under different scenarios. Specifically, we will conduct a series of in-depth evaluations (i.e., Exp-3 to Exp-7) using GPT-4o.
Overall Results of GTP-4o.
The overall result of GPT-4o is shown in Figure 5. These heatmaps visualize the performance of GPT-4o on various low-level ChartQA tasks under different prompt conditions. The progression from subfigures (a) to (d) clearly indicates the incremental benefits of incorporating visual prompts, optimization strategies, and their combination, culminating in the most effective approach for improving GPT-4o’s performance in low-level ChartQA tasks.
Figure 4: Effectiveness of GPT-4o
Figure 5: The Effectiveness of GPT-4o across 10 Low-level Tasks and 4 Question Types
@@ 3.2 Experimental Results and Findings
As depicted in the bar chart at the top of Figure 4, GPT-4o’s accuracy reaches near 90% for basic bar charts, but hovers around 66% for similar tasks involving other chart types. This performance gap suggests that, despite the dataset containing a significant number of reasoning tasks that are generally straightforward for humans, the effectiveness of GPT-4o in ChartQA tasks has not yet reached that of the average human.
Finding-4: The performance of GPT-4o declines as task complexity increases, mirroring human performance, but it has not yet matched the analytical capabilities of average humans.
◆ Exp-2: Impact of Question Types
The experimental settings are the same as in Exp-1.
The Effectiveness of Question Types.
Figure 5 (a) provides an overview of GPT-4o’s performance across 10 low-level tasks using four different textual prompts. Notably, GPT-4o achieves the highest overall accuracy of 77.8% in “Multiple-Choice” questions. It also demonstrates a strong performance in “Yes or No” questions, with an accuracy of 74.8%. Comparatively, GPT-4o performs relatively better in “Multiple-Choice” and “Yes-or-No” questions compared to “Fill-in-the-Blank” and “Error Correction” questions. The former prompt types inherently provide GPT-4o with answer options to select or evaluate, while the latter two require direct answer generation from GPT-4o.
@@
Finding-5: Structured textual prompts and candidate answers significantly enhance GPT-4o’s ability to reason out correct responses.
◆ Exp-3: Optimization of Textual Prompts
This experiment will examine the influence of commonly used textual prompt optimization strategies on MLLMs, including RolePlay Shanahan et al. (2023), Tutorial, and ChartCoT Xu et al. (2023), which is based on the Chain-of-Thought (CoT) prompt strategy Wei et al. (2022). The Chain-of-Thought (CoT) prompt strategy has proven effective in various scenarios Wei et al. (2022). CoT aims to guide the model by mimicking the step-by-step reasoning process humans use to solve problems. Recently, Xu et al. Xu et al. (2023) implemented the CoT strategy for ChartQA tasks, namely ChartCoT Xu et al. (2023). ChartCoT poses a series of questions to guide the model in understanding the chart’s details before formulating an answer. However, ChartCoT faces challenges in ensuring the accuracy of GPT-4o’s responses to guiding questions, especially with complex charts.
Chain-of-Charts Prompts.
Therefore, we introduce a novel prompting strategy, termed Chain-of-Charts, which builds based on the Chain-of-Thoughts approach Wei et al. (2022), as shown in Appendix F.6. The core of Chain-of-Charts lies in orchestrating a sequence of questions and their corresponding answers (($q_{1}$, $a_{1}$), ($q_{2}$, $a_{2}$),…($q_{m}$, $a_{m}$)) to progressively guide the model towards a deeper understanding of the chart’s details, thereby enhancing its ability to formulate accurate responses.
Experimental Settings.
Shanahan et al. Shanahan et al. (2023) suggested that giving large models specific roles could enhance their performance on particular tasks. Inspired by this, we assigned GPT-4o the role of a visualization expert to see if it would enhance its performance. Our observations showed that more detailed prompts resulted in more precise and accurate responses from GPT-4o. As a result, we created a detailed ChartQA tutorial called the Tutorial prompt. Examples of these two prompts can be found in Appendix F.3 and Appendix F.4 respectively.
Overall Results.
Figure 5(b) reports the performance of GPT-4o with Chain-of-Charts. Compared with Figure 5(a), we can observe a significant enhancement in GPT-4o’s capabilities across 10 tasks and four question types. Table 4(a) shows the overall accuracy of GPT-4o on 10 low-level tasks under 5 Textual Prompt strategies. Overall, Chain-of-Charts leads in average accuracy across all tasks with 83.5%, outperforming ChartCoT’s accuracy of 76.1% by 6.9%. Specifically, Chain-of-Charts achieves the highest accuracy on five tasks including Reasoning, Determine Range, Order, Filter, Retrieve Value, Find Extremumm and Find Cluster, with accuracy of 78.9%, 89.2%, 86.3%, 76.7%, 96.7%, 97.5% and 92.5%. These tasks demand precise reasoning from GPT-4o, based on accurate identification of element coordinates and values. The Chain-of-Charts prompt framework effectively provides GPT-4o with the correct value and coordinate references, significantly aiding in the accurate positioning of different elements.
Finding-6: Chain-of-Charts supplies GPT-4o with accurate chart reference information, enhancing the model’s comprehension and detailed reasoning of chart structures and elements.
◆ Exp-4: Impact of Visual Prompts
Experimental Settings.
@@
In this experiment, we design three types of visual prompts based on graphical overlay strategies Kong and Agrawala (2012), namely, handwriting, regular shape, and special design, as shown in Figure 9. The design of visual prompts are detailed in Appendix F.2. Specifically, we generate 255 visual prompts for 35 charts as the testing samples. These 255 visual prompts are associated with 1,020 test samples and cover 10 low-level tasks. Please refer to Step-5 in the Appendix A.2 for details.
Overall Results.
@@
Figure 5(c) demonstrates the strong performance of GPT-4o with visual prompts across 10 tasks and 4 question types. Additionally, Figure 22(a)-(c) showcases the performance of visual prompts under different textual prompts, low-level ChartQA tasks, and chart types. In general, visual prompts enhance GPT-4o’s performance. Particularly, Figure  22(b) reveals significant improvements in Reasoning and Anomaly Detection tasks, indicating the model’s ability to accurately analyze and reason with relevant data.
Finding-7: Visual prompts greatly improve GPT-4o’s performance, showing the value of visual information for comprehension and reasoning.